OpenStudy (anonymous):

A 0.40 kg rope is stretched between two supports 4.8 m apart. When one support is struck by a hammer, a transverse wave travels down the rope and reaches the other support in 0.85 s. What is the tension in the rope?

OpenStudy (anonymous):

What you need to do is to draw a free-body diagram first...Hmmm...I had my mechanics two semesters ago and I kinda need to refresh on this one first... ^_^

OpenStudy (anonymous):

The wave velocity(v) and the Tension(T) in the rope for transverse waves are related as \[v=\sqrt{\frac{T}{\mu}}\]. For more on how you get it refer http://www.yosunism.com/undergradphysics/transversewaves.pdf Here $$ v=\frac{4.8}{0.85}=5.64ms^{-1} $$ and $$ \mu =\frac{0.40}{4.8}=0.08 kgm^{-1}$$ Plugging these in will give you $$T=\mu v^2 =2.66N $$